# GAEA: A Geolocation Aware Conversational Model

### v2

@@ Abstract
Image geolocalization, in which, traditionally, an AI model predicts the precise GPS coordinates of an image is a challenging task with many downstream applications. However, the user cannot utilize the model to further their knowledge other than the GPS coordinate; the model lacks an understanding of the location and the conversational ability to communicate with the user. In recent days, with tremendous progress of large multimodal models (LMMs)—proprietary and open-source—researchers have attempted to geolocalize images via LMMs. However, the issues remain unaddressed; beyond general tasks, for more specialized downstream tasks, one of which is geolocalization, LMMs struggle. In this work, we propose to solve this problem by introducing a conversational model GAEA that can provide information regarding the location of an image, as required by a user. No large-scale dataset enabling the training of such a model exists. Thus we propose GAEA-1.6M, a comprehensive dataset with 800K images and around 1.6M question-answer pairs constructed by leveraging OpenStreetMap (OSM) attributes and geographical context clues. For quantitative evaluation, we propose a diverse benchmark, GAEA-Bench, comprising 4K image-text pairs to evaluate conversational capabilities equipped with diverse question types. We consider 11 state-of-the-art open-source and proprietary LMMs and demonstrate that GAEA significantly outperforms the best open-source model, LLaVA-OneVision by 25.69% and the best proprietary model, GPT-4o by 8.28%. Our dataset, model and codes are available https://ucf-crcv.github.io/GAEA.
Figure 1: Performance of GAEA and other LMMs on global scale image geo-localization. GAEA makes correct predictions when asked different questions about summarizing a scene, location, and geographical context. While GPT-4o-mini can give correct suggestions relevant to the region, GAEA provides a correct amenity with proximity to the location of the image.
@@ 1 Introduction
Figure 2: Data Collection and Annotation Pipeline. GAEA-1.6M includes geographically diverse visual samples from various data sources, such as MP-16, GLD-v2, and CityGuesser68k (left). We also incorporate OpenStreetMap (OSM) metadata and auxiliary context for each image, ranging from climate zones to geographical clues about the country (middle). Using open-source LLMs and GPT-4o, we generate four diverse question-answer pairs across geolocation, reasoning, and conversational subsets (right).
@@
Image geolocalization [59, 50, 23, 32, 55] is a notoriously challenging task, in which, conventional AI models predict the precise GPS coordinate of an image taken anywhere on Earth. Moreover, seasonal changes, geographical and climatic diversity, changes in solar zenith angle, and lack of diverse image distributions make the geolocalization task more challenging. Although difficult, geolocalization has direct applications in multiple domains, including tourism, navigation [16], urban planning [46], and security [50], among many.
@@
Recently, CLIP-inspired image-to-GPS retrieval approach, GeoCLIP [50], has shown significant performance in global-scale image geolocalization. To further mitigate the performance gap, and to increase the generalization capacity of the models, interestingly, a new wave of works infuse human-level cognition and inference capacity in their model training [23, 32, 55]. E.g., PIGEON is trained on data from the popular geolocalization game GeoGuesser [2]; a recent vision-language model, GeoReasoner [32] use user- and administrator-maintained approximately 3K textual clues from GeoGuessr and Tuxun gaming platforms.
Figure 3: Overview of GAEA-Bench. GAEA-Bench is designed to evaluate the conversational abilities of various LMMs across different question types, including MCQs, T/F, and both short and long VQAs. We have carefully selected a subset of 4k samples from MP-16 and generated corresponding OSM metadata to generate QA pairs using GPT-4o. GAEA-Bench aims to fill the gap in conversational benchmarks by incorporating geolocalization capabilities.
@@
These focused geolocalization models, however, lack geographical understanding of the predicted locations beyond their GPS coordinates. They cannot provide additional information that might be invaluable for applications such as tourism, navigation, urban planning, etc. Even if the models possess that understanding, they do not have the conversational ability to convey that information and fail to meet the user’s needs. In contrast, despite having the conversational capability, visually and textually prompted large language models (LLMs) [20, 56, 48] and their multimodal variants, popularly referred to as large multimodal models (LMMs) [14, 10, 47, 11, 35], fail to capture fine-grained nuances from an image in specialized downstream tasks such as geolocalization, making their predictions vastly imprecise and worse than random guesses in many cases; see Figure 1.
@@
Motivated by all these aspects, in this paper, we propose GAEA, an open-source conversational model with a global-scale geolocalization capability. To the best of our knowledge, this is the first work in the ground-view geolocalization domain that introduces an open-source conversational chatbot, where the user can obtain image geolocalization, relevant description of the image, and engage in a meaningful conversation about the surrounding landmarks, natural attractions, restaurants or coffee shops, medical or emergency facilities, and recreational areas.
@@
However, training an open-source LMM with conversational capacity is not straightforward. These models are data-hungry and their training is compute intensive. Unfortunately, no dataset can facilitate the training of such a model. To this end, we meticulously curate a GAEA-1.6M —a high-quality conversational VQA pair equipped with diversity in scene understanding and image captions for training and instruction tuning the LMMs on the street-level geolocalization task. GAEA-1.6M is a comprehensive dataset consisting over 800K images from MP-16 [29], covering locations around the Earth. We augment these images with rich meta-data from the OpenStreetMap (OSM) [38] at a 1km radius, a first effort of its kind. OSM attributes contain details about the surrounding area, nearby landmarks, accessible services, and historical buildup of the region. The QA subset of GAEA-1.6M contains 380K QA pairs; the geolocalizable explanatory captions set contains 385K images and is equipped with their corresponding knowledge and reasoning captions. These knowledge and reasoning captions are constructed using a set of geographical context clues from GeoGuessr [2] that enable the model to gain a holistic understanding of the location. Taken together, GAEA-1.6M is the largest and most comprehensive collection of geolocalizable and conversational QA pairs. We use this data source to design our conversational chatbot, GAEA.
@@
To quantitatively evaluate the conversational capability of LMMs and address the scarcity of benchmark datasets in a geolocalization setting, we propose GAEA-Bench, a diverse set of 4K conversational question samples.
@@
  • We propose GAEA-1.6M (§Section 3), a new dataset for training conversational image geolocalization models.
  • For evaluating conversational capabilities in geolocalization setting (§Section 5), we propose GAEA-Bench, a novel benchmark with different types of question-answers.
  • We propose GAEA, a conversational chatbot (§4) that goes beyond global-scale geolocalization and provides information about the location described by an image.
  • We quantitatively compare the performance of our model to 8 state-of-the-art open-source, and 3 proprietary LMMs, including GPT-4o [9] and Gemini-2.0-Flash [47].
@@ 2 Related Work
Vision language models (VLMs) have been at the forefront of computer vision research; geo-localizable VLMs are in their nascent stages.
@@
Vision Language Models. Multimodal learning unifies different modalities by a common representation. By contrastively fitting text and images into the same feature space, CLIP [39] has revolutionized multimodal learning. LLMs like GPT2 [40] made strides in representing text and next token prediction. Visual question answering (VQA) was of interest before, but, after LLaVA [34] and BLIP2 [31] combined the conversational aspects of LLMs and the representational capabilities of CLIP-like models, many problems of VQA are addressed.After that, numerous modern works emerged, such as GeoChat [27], Qwen2.5-VL [12], LLaMA-3.2 Vision [10], and LLaVA-OneVision [30] as well as proprietary models like GPT4 [9] and Gemini [47]. Even though most of these models are excellent for general VQA, they perform poorly on specialized tasks in fields like medicine, statistics, and geolocalization. This inspires the need for specialized VLMs that can address specific tasks.
Figure 4: Qualitative examples showcasing various question-types, including multiple-choice, true/false, short and long VQAs generated using an open-source model on our GAEA-1.6M dataset. We carefully select geographical tags from OSM metadata to generate QA pairs.
@@
Geo-localization is a crucial field in vision research with essential applications in forensics, social media, and exploration; see [16, 46, 50]. On a global scale, Weyand et al. [53] first introduced a classification-based approach on the Im2GPS [24] dataset. Vo et al. [51] introduced classification in multiple hierarchies, while CPlaNet [45] introduced a combinatorial partitioning technique for combining coarse hierarchies to predict finer ones. Over the years many other works like ISNs [25], TransGeo [59], TransLocator [52], and GeoDecoder [15] have made significant advancements in this classification-based worldwide geolocalization by introducing scene-based specialized encoders and hierarchical evaluation, auxiliary scene recognition, and twin encoder approach, and query-based encoder-decoder architecture respectively. PIGEON [23], the most recent work, leverages the image representation capabilities of the CLIP vision encoder, and a unique clustering approach to improve its geo-localization performance. Image-to-image retrieval models tend to be more accurate than their classification-based counterparts but infeasible on a global scale due to their requirement for large reference image galleries. GeoCLIP [50] was the first work to incorporate the contrastive multimodal learning between images and raw GPS information that revolutionized this domain by introducing a more accurate retrieval-based model for a global scale.
@@
These specialized models work well for worldwide image geo-localization but lack the conversational aspect that can aid an individual in gaining a holistic understanding of a location portrayed in an image. GeoReasoner [32] attempts to incorporate an inherent geospatial understanding into a VLM by looking at specific information displayed in the image. It also introduces the idea of locatability, which can determine the extent of that information present in the image which may improve the reasoning capability of the model. The model however lacks the conversational aspect, and the locatability-based filtering of data might hurt its generalization capability. We address these issues in GAEA by primarily focusing on its conversational ability.
@@
The generalizability of GAEA comes from its training data. All specialized geo-localization models that function on a global scale train their model on MP-16 [29] which is a large-scale worldwide dataset. However, it lacks the verbal context required in VLM training. Hence, we introduce a new conversational dataset GAEA-1.6M; see details in §Section 3. Additionally, we introduce the first conversational benchmark in §Section 4 to evaluate Geolocalization VLMs and an evaluation pipeline to judge the efficacy of such models.
3 GAEA-1.6M
@@
The GAEA-1.6M dataset provides comprehensive global coverage, featuring both rich conversational and diverse geolocalization sets. It includes various QA formats, such as MCQs, True/False, and open-ended VQA (long and short), from more than 234 countries/territories, grouped under conversational and geolocalization groups. Spanning 40k cities across 7 continents, GAEA-1.6M is structured into two key groups: conversational and geolocalization. With over 1.6 million QA pairs, it captures the geographical diversity of both underrepresented and widely recognized regions worldwide.
3.1 Dataset Curation and Annotation
@@
Acquiring Diverse Geo-localizable Images. We sample geographically diverse visual data from MediaEval 2016 (MP-16) [29], Google Landmarks v2 (GLDv2) [54], and CityGuessr [28] to curate GAEA-1.6M.
@@
To achieve a balanced geographical distribution in GAEA-1.6M, we use the 10th hierarchy of S2-Cells [5] to partition our filtered MP-16 dataset into 16,753 spatial grid cells. S2-Cells enable hierarchical spatial indexing, ensuring diverse global coverage while preventing overrepresenting densely imaged regions. We randomly sample up to 200 images from each cell, resulting in a final set of over 750K distinct samples.
@@
CityGuessr68k [28] focuses on global video-based geolocalization emphasizing urban regions and hierarchical prediction across 166 major cities. To incorporate this diversity, we randomly sample one frame from each of the 54K training videos and include them in our dataset. These three sources provide over 852K geographically diverse geolocalizable images, forming GAEA-1.6M dataset.
Figure 5: Evaluation pipeline highlights various question types we introduce in our GAEA-1.6M. We use GPT-4o as a judge to score such responses.
3.2 Meta-data curation for dataset annotation
@@
OSM data is multilingual, which is a key challenge. To ensure accessibility, we use GPT-4o [9] to translate these annotations into English. Additionally, many retrieved tags consisted of plain numbers or non-meaningful entries, which we systematically filtered out to retain only informative and contextually relevant metadata. To our knowledge, this is the first work to utilize OSM’s rich metadata to develop a conversational chatbot for ground-view geolocalization. We curated the OSM metadata for MP-16 and GLDv2 visual samples.
Curating Country-Specific Geographical Clues. We web-crawled diverse clues from Plonkit[4], an open-source community resource for the GeoGuessr [2] game, which has over 65 million players. Similar datasets have been used in recent works [23, 32]. We obtained 129 country clues but found gaps for some countries, such as New Zealand and France. To address this, we curated clues for 58 additional countries using GPT-4o, aligning them with Plonkit’s style, resulting in altogether 187 countries. These clues are incorporated into our dataset for generating reasoning-based QAs. For examples of the type of clues utilized, see Figure 19 in the Appendix.
@@
3.3 Question-Answer (QA) Pairs Generation
As seen in Figure 2, GAEA-1.6M is carefully curated to enhance ground-view geolocalization through diverse, context-rich QA pairs. Comprising over 800K distinct images and around 1.6 million QA pairs, it stands as the largest and most comprehensive dataset for this task; see Figure 4. Unlike existing works, such as [32, 19], which are limited to JSON structures and fewer question types, our work emphasizes the conversational capabilities of the model, providing a broader range of QA formats. The dataset is divided into three subsets—Conversational, Reasoning, and Geo-Localization, each designed to capture different aspects of geographic understanding. These subsets feature various question formats, including multiple-choice [37], true/false, and open-ended questions (SVQA and LVQA) [49]. Below, we detail the curation process for each subset.
Conversational QA Generation. We generate conversational QA pairs using OSM metadata from the sampled MP-16 and GLDv2 subsets. We prompt Qwen-2.5-14B [56] with enriched OSM attributes to create diverse question formats, including short-form, multiple-choice, and true/false questions. These OSM tags cover various categories such as amenities, food places, financial institutions, government offices, accommodation, transportation, healthcare, religious sites, education, and waterways. This subset comprises over 380K questions.

### v3

# GAEA: A Geolocation Aware Conversational Assistant

@@ Abstract
Image geolocalization, in which an AI model traditionally predicts the precise GPS coordinates of an image, is a challenging task with many downstream applications. However, the user cannot utilize the model to further their knowledge beyond the GPS coordinates; the model lacks an understanding of the location and the conversational ability to communicate with the user. In recent days, with the tremendous progress of large multimodal models (LMMs)—proprietary and open-source—researchers have attempted to geolocalize images via LMMs. However, the issues remain unaddressed; beyond general tasks, for more specialized downstream tasks, such as geolocalization, LMMs struggle. In this work, we propose solving this problem by introducing a conversational model, GAEA, that provides information regarding the location of an image as the user requires. No large-scale dataset enabling the training of such a model exists. Thus, we propose GAEA-1.4M, a comprehensive dataset comprising over 800k images and approximately 1.4M question-answer pairs, constructed by leveraging OpenStreetMap (OSM) attributes and geographical context clues. For quantitative evaluation, we propose a diverse benchmark, GAEA-Bench, comprising 3.5k image-text pairs to evaluate conversational capabilities equipped with diverse question types. We consider 11 state-of-the-art open-source and proprietary LMMs and demonstrate that GAEA significantly outperforms the best open-source model, LLaVA-OneVision, by 18.2% and the best proprietary model, GPT-4o, by 7.2%. Our dataset, model and codes are available https://ucf-crcv.github.io/GAEA.
Figure 1: We compare the performance of various LMMs on the geographically‑grounded visual‑question‑answering task, included in our new GAEA-Bench benchmark. Most LMMs can describe the Wat Pho statue, but only GAEA, our Geolocation AwarE Assistant, retrieves the correct nearby cafe, Cafe Amazon.
@@ 1 Introduction
Image geolocalization [63, 52, 25, 34, 58] is a notoriously challenging task, in which, conventional AI models predict the precise GPS coordinate of an image taken anywhere on Earth. Moreover, seasonal changes, geographical and climatic diversity, changes in solar zenith angle, and lack of diverse image distributions make the geolocalization task more challenging. Although challenging, geolocalization has direct applications in multiple domains, including tourism [16], navigation, urban planning [48], and security [52, 22], among others.
@@
Recently, the CLIP-inspired image-to-GPS retrieval approach, GeoCLIP [52], has shown significant performance in global-scale image geolocalization. To further mitigate the performance gap, and to increase the generalization capacity of the models, interestingly, a new wave of works infuse human-level cognition and inference capacity in their model training [25, 34, 58]. E.g., PIGEON is trained on data from the popular geolocalization game GeoGuesser [2]; a recent vision-language model, GeoReasoner [34] uses user- and administrator-maintained approximately 3K textual clues from GeoGuessr and Tuxun gaming platforms.
@@
These focused geolocalization models lack a geographical understanding of the predicted locations beyond their GPS coordinates. They cannot provide additional information that might be invaluable for applications such as tourism, navigation, urban planning, etc. Even if the models possess that understanding, they do not possess the conversational ability to convey that information and fail to meet the user’s needs. In contrast, despite having the conversational capability, visually and textually prompted large language models (LLMs) [21, 59, 50] and their multimodal variants, popularly referred to as large multimodal models (LMMs) [14, 10, 49, 11, 37], fail to capture fine-grained nuances from an image in specialized downstream tasks such as geolocalization, making their predictions vastly imprecise and worse than random guesses in many cases; see Figure 1.
Figure 2: Data Collection and Annotation Pipeline. GAEA-1.4M includes geographically diverse visual samples from various data sources, such as MP-16 [31], GLD-v2 [57], and CityGuesser68k [30] (left). We also incorporate additional metadata and auxiliary context for each image from OpenStreetMap (OSM), ranging from climate zones to geographical clues about the country (middle). Using open-source LLMs and GPT-4o, we generate four diverse question-answer pairs across geolocation, reasoning, and conversational subsets (right).
@@
Motivated by these aspects, in this paper, we propose GAEA, an open-source conversational model with global-scale geolocalization capability. To the best of our knowledge, this is the first work in the ground-view geolocalization domain that introduces an open-source conversational chatbot, where the user can obtain image geolocalization, relevant description of the image, and engage in a meaningful conversation about the surrounding landmarks, natural attractions, restaurants or coffee shops, medical or emergency facilities, and recreational areas.
@@
However, training an open-source LMM with conversational capacity is not straightforward. These models are data-hungry, and their training is computationally intensive. Unfortunately, no dataset can facilitate the training of such a model. To this end, we meticulously curate a GAEA-1.4M —a high-quality conversational VQA pair equipped with diversity in scene understanding and image captions for training and instruction tuning the LMMs on the street-level geolocalization task. GAEA-1.4M is a comprehensive dataset consisting over 800k images from MP-16 [31], GLD-v2 [57], and CityGuessr68k [30] covering locations around the Earth. We augment these images with rich metadata from the OpenStreetMap (OSM) [40] at a 1 km radius, a first effort of its kind. OSM attributes contain details about the surrounding area, nearby landmarks, accessible services, and historical buildup of the region. Additionally, the geolocalizable explanatory captions set contains 385K image-QA pairs and is equipped with their corresponding knowledge and reasoning captions. These knowledge and reasoning captions are constructed using a set of geographical context clues from GeoGuessr [2] that enable the model to gain a holistic understanding of the location. Taken together, GAEA-1.4M is the largest and most comprehensive collection of geolocalizable and conversational QA pairs. Using this, we train our conversational chatbot, GAEA.
@@
To quantitatively evaluate the conversational capability of LMMs and address the scarcity of benchmark datasets in a geolocalization setting, we propose GAEA-Bench, a diverse set of 3.5k conversational question samples.
@@
  • We propose GAEA-1.4M (§Section 3), a new dataset for training conversational image geolocalization models.
  • For evaluating conversational capabilities in a geolocalization setting (§Section 5), we propose GAEA-Bench, a novel benchmark of 3.5k samples with various question types, including short, long, MCQs, and T/F.
  • We propose GAEA, a conversational chatbot (§4) that extends beyond global-scale geolocalization, providing information about the location described by an image.
  • We quantitatively compare the performance of our model to 8 state-of-the-art open-source and 3 proprietary LMMs, including GPT-4o [9] and Gemini-2.0-Flash [49].
@@ 2 Related Work
Large Multimodal Models (LMMs) have been at the forefront of computer vision research; geo-localizable LMMs are in their nascent stages. Multimodal learning unifies different modalities by a common representation. By contrastively fitting text and images into the same feature space, CLIP [41] has revolutionized multimodal learning. LLMs like GPT2 [42] made strides in representing text and next token prediction. Visual question answering (VQA) was of interest before, but after LLaVA [36] and BLIP2 [33] combined the conversational aspects of LLMs and the representational capabilities of CLIP-like models, many problems of VQA are addressed. After that, numerous modern works emerged, such as GeoChat [29], Qwen2.5-VL [12], LLaMA-3.2 Vision [10], and LLaVA-OneVision [32] as well as proprietary models like GPT4 [9] and Gemini [49]. Although most of these models are excellent for general VQA, they perform poorly on specialized tasks in fields like medicine, statistics, and geolocalization. This inspires the need for specialized LMMs that can address specific tasks.
@@
Geo-localization is a crucial field in vision research with essential applications in forensics, social media, and exploration; see [16, 48, 52]. On a global scale, Weyand et al. [56] first introduced a classification-based approach on the Im2GPS [26] dataset. Vo et al. [53] introduced classification in multiple hierarchies, while CPlaNet [47] introduced a combinatorial partitioning technique for combining coarse hierarchies to predict finer ones. Over the years many other works like ISNs [27], TransGeo [63], TransLocator [54], and GeoDecoder [15] have made significant advancements in this classification-based worldwide geolocalization by introducing scene-based specialized encoders and hierarchical evaluation, auxiliary scene recognition, and twin encoder approach, and a query-based encoder-decoder architecture, respectively. PIGEON [25], the most recent work, leverages the image representation capabilities of the CLIP vision encoder and a unique clustering approach to improve its geo-localization performance. Image-to-image retrieval models tend to be more accurate than their classification-based counterparts, but are infeasible on a global scale due to their requirement for large reference image galleries. GeoCLIP [52] was the first work to incorporate the contrastive multimodal learning between images and raw GPS information that revolutionized this domain by introducing a more accurate retrieval-based model for a global scale.
@@
These specialized models work well for worldwide image geo-localization but lack the conversational aspect that can aid an individual in gaining a holistic understanding of a location portrayed in an image. GeoReasoner [34] attempts to incorporate an inherent geospatial understanding into a LMM by looking at specific information displayed in the image. It also introduces the idea of locatability, which can determine the extent of that information present in the image, which may improve the reasoning capability of the model. The model, however, lacks the conversational aspect, and the locatability-based filtering of data might hurt its generalization capability. We address these issues in GAEA by primarily focusing on its conversational ability.
@@
The generalizability of GAEA comes from its training data. All specialized geo-localization models that function on a global scale train their model on MP-16 [31], which is a large-scale worldwide dataset. However, it lacks the verbal context required in LMM training. Hence, we introduce a new conversational dataset GAEA-1.4M; see details in §Section 3. Additionally, we introduce the first conversational benchmark in §Section 4 to evaluate Geolocalization LMMs and an evaluation pipeline to judge the efficacy of such models.
Figure 3: Qualitative samples from our GAEA-1.4M training set showcasing various question-types, including multiple-choice, true/false, short and long VQAs generated using a proprietary model, GPT-4o [39]. We carefully select geographical tags from OSM metadata to generate conversational question-answer pairs.
3 Dataset Curation and Annotation
3.1 GAEA-1.4M
@@ 3 Dataset Curation and Annotation
The GAEA-1.4M dataset provides comprehensive global coverage, featuring both rich conversational and diverse geolocalization sets. It includes various QA formats, such as MCQs, True/False, and open-ended VQA (long and short), from more than 234 countries/territories, grouped under conversational and geolocalization groups. Spanning 40k cities across 7 continents, GAEA-1.4M is structured into two key groups: conversational and geolocalization. With over 1.4 million QA pairs, it captures the geographical diversity of both underrepresented and widely recognized regions worldwide. Figure 2 shows our complete curation pipeline, which we will discuss piecewise.
@@
Acquiring Diverse Geo-localizable Images. We sample geographically diverse visual data from MediaEval 2016 (MP-16) [31], Google Landmarks v2 (GLDv2) [57], and CityGuessr [30] to curate GAEA-1.4M.
@@
To achieve a balanced geographical distribution in GAEA-1.4M, we use the 10th hierarchy of S2-Cells [5] to partition our filtered MP-16 dataset into 16,753 spatial grid cells. S2-Cells enable hierarchical spatial indexing, ensuring diverse global coverage while preventing overrepresenting densely imaged regions. We randomly sample up to 200 images from each cell, resulting in a final set of over 750k distinct samples.
@@
CityGuessr68k [30] focuses on global video-based geolocalization emphasizing urban regions and hierarchical prediction across 166 major cities. To incorporate this diversity, we randomly sample one frame from each of the 54k training videos and include them in our dataset. These three sources provide over 852k geographically diverse geolocalizable images, forming GAEA-1.4M dataset.
3.1.1 Meta-data curation for dataset annotation
@@
OSM data is multilingual, which is a key challenge. To ensure accessibility, we use GPT-4o [9] to translate these annotations into English. Additionally, many retrieved tags consisted of plain numbers or non-meaningful entries, which we systematically filtered out to retain only informative and contextually relevant metadata. To our knowledge, this is the first work to utilize OSM’s rich metadata to develop a conversational chatbot for ground-view geolocalization. Figure 3 shows high-priority meta-tags derived from OSM churned data.
Curating Country-Specific Geographical Clues. We web-crawled diverse clues from Plonkit[4], an open-source community resource for the GeoGuessr [2] game, which has over 65 million players. Similar datasets have been used in recent works [25, 34]. We obtained 129 country clues but found gaps for some countries, such as New Zealand and France. To address this, we curated clues for 58 additional countries using GPT-4o, aligning them with Plonkit’s style, resulting in altogether 187 countries. These clues are incorporated into our dataset for generating reasoning-based QAs. For examples of the type of clues utilized, see Figure 9 in the Supplementary material.
@@
Figure 4: Overview of GAEA-Bench. GAEA-Bench is designed to evaluate the conversational abilities of various LMMs across different question types, including MCQs, T/F, and both short and long VQAs. We have carefully selected a subset of 3.5k samples from MP-16 [31] and generated corresponding OSM metadata to generate QA pairs using GPT-4o [39]. GAEA-Bench aims to fill the gap in conversational benchmarks by incorporating geolocalization capabilities.
3.1.2 Question-Answer (QA) Pairs Generation
GAEA-1.4M is carefully curated to enhance ground-view geolocalization through diverse, context-rich QA pairs; see Figure 2. Comprising over 800k distinct images and around 1.4 million QA pairs, it stands as the largest and most comprehensive dataset for this task; see Figure 3. Unlike existing works, such as [34, 20], which are limited to JSON structures and fewer question types, our work emphasizes the conversational capabilities of the model, providing a broader range of QA formats. The dataset is divided into three subsets—Conversational, Reasoning, and Geo-Localization, each designed to capture different aspects of geographic understanding. These subsets feature various question formats, including multiple-choice [38], true/false, and open-ended questions (SVQA and LVQA) [51]. Below, we detail the curation process for each subset.
Conversational QA Generation. We generate conversational QA pairs using OSM metadata from the sampled MP-16 and GLDv2 subsets. We prompt Qwen-2.5-14B [59] with enriched OSM attributes to create diverse question formats, including short-form, multiple-choice, and true/false questions. These OSM tags cover various categories such as amenities, food places, financial institutions, government offices, accommodation, transportation, healthcare, religious sites, education, and waterways. This subset comprises over 380k questions. Figure 3 showcases qualitative samples of QA pairs curated using OSM churned data.